If the sum of the zeroes of the polynomial kx²+2kx+4k is equal to twice Thier product find the value of k

Grade:10

2 Answers

V RAHUL
6 years ago
Keep can have Any value as the roots are not dependent on the values of k
If we are going to take common from the quadratic equation here
Then we can observe that the value of k is not effecting the resultant roots here therefore 
K can have any numerical value possible
Harshit
6 years ago
sum of roots=2(product of roots)     -2k/k=2(4k/k)   -2=8 It is not possible . so their is no real of k for which this statement is ture